# NeRC: Neural Ranging Correction through Differentiable Moving Horizon Location Estimation

### v1

@@ Acknowledgements.
- Wang et al. (2015b) Yang Wang, Rong Yang, Keck Voon Ling, and Eng Kee Poh. 2015b. Robust vector tracking loop using moving horizon estimation. In Proceedings of the ION 2015 Pacific PNT Meeting. 640–648.
- Wen et al. (2021) Weisong Wen, Tim Pfeifer, Xiwei Bai, and Li-Ta Hsu. 2021. Factor graph optimization for GNSS/INS integration: A comparison with the extended Kalman filter. NAVIGATION: Journal of the Institute of Navigation 68, 2 (2021), 315–331.
- Wen et al. (2019) Weisong Wen, Guohao Zhang, and Li-ta Hsu. 2019. Correcting NLOS by 3D LiDAR and building height to improve GNSS single point positioning. Navigation 66, 4 (2019), 705–718.
- Weng et al. (2023) Duojie Weng, Zhiyu Hou, Yang Meng, Miaomiao Cai, and Yanyiu Chan. 2023. Characterization and mitigation of urban GNSS multipath effects on smartphones. Measurement 223 (2023), 113766.
- Weng et al. (2024a) Xu Weng, Yuhui Jin, and KV Ling. 2024a. GnssQuest: Questing for Suitable GNSS Satellites through Augmented Reality. In Proceedings of the 22nd ACM Conference on Embedded Networked Sensor Systems. 867–868.
- Weng et al. (2024c) Xu Weng, K.V. Ling, Haochen Liu, and Kun Cao. 2024c. Towards End-to-End GPS Localization with Neural Pseudorange Correction. In 2024 27th International Conference on Information Fusion (FUSION). 1–7. doi:10.23919/FUSION59988.2024.10706359
- Weng and Ling (2023) Xu Weng and Keck Voon Ling. 2023. Localization with noisy Android raw GNSS measurements. In 2023 IEEE Asia Pacific Conference on Wireless and Mobile (APWiMob). IEEE, 95–101.
@@
- Xia et al. (2024) Xiao Xia, Weisong Wen, and Li-Ta Hsu. 2024. Integrity-Constrained Factor Graph Optimization for GNSS Positioning in Urban Canyons. NAVIGATION: Journal of the Institute of Navigation 71, 3 (2024).
@@
- Xu et al. (2023b) Huatao Xu, Pengfei Zhou, Rui Tan, and Mo Li. 2023b. Practically adopting human activity recognition. In Proceedings of the 29th Annual International Conference on Mobile Computing and Networking. 1–15.
- Xu and Hsu (2024) Penghui Xu and Li-Ta Hsu. 2024. AutoW: Self-Supervision Learning for Weighting Estimation in GNSS Positioning. In Proceedings of the 37th International Technical Meeting of the Satellite Division of The Institute of Navigation (ION GNSS+ 2024). 2630–2644.
- Xu et al. (2023a) Penghui Xu, Hoi-Fung Ng, Yihan Zhong, Guohao Zhang, Weisong Wen, Bo Yang, and Li-Ta Hsu. 2023a. Differentiable factor graph optimization with intelligent covariance adaptation for accurate smartphone positioning. In Proceedings of the 36th International Technical Meeting of the Satellite Division of The Institute of Navigation (ION GNSS+ 2023). 2765–2773.
@@
- Yang et al. (2023a) Fan Yang, Chen Wang, Cesar Cadena, and Marco Hutter. 2023a. iPlanner: Imperative Path Planning. In Robotics: Science and Systems (RSS). https://arxiv.org/abs/2302.11434
- Yang et al. (2023b) Fan Yang, Chen Wang, Cesar Cadena, and Marco Hutter. 2023b. iPlanner: Imperative Path Planning. In Proceedings of Robotics: Science and Systems. Daegu, Republic of Korea. doi:10.15607/RSS.2023.XIX.064
- Zhan et al. (2024) Zitong Zhan, Dasong Gao, Yun-Jou Lin, Youjie Xia, and Chen Wang. 2024. iMatching: Imperative correspondence learning. In European Conference on Computer Vision. Springer, 183–200.
- Zhang et al. (2023) Aston Zhang, Zachary C Lipton, Mu Li, and Alexander J Smola. 2023. Dive into deep learning. Cambridge University Press.
@@
- Zhao et al. (2025) Ling Zhao, Zhengliang Lu, Keck Voon Ling, Yuandong Hu, Kan Zheng, and Wenhe Liao. 2025. Multi-rate Cascade Spacecraft Attitude-Orbit Integrated State Estimation and Control Framework based on MHE and PWA-MPC. IEEE Trans. Aerospace Electron. Systems (2025).
- Zhao et al. (2023) Xiaopeng Zhao, Zhenlin An, Qingrui Pan, and Lei Yang. 2023. Nerf2: Neural radio-frequency radiance fields. In Proceedings of the 29th Annual International Conference on Mobile Computing and Networking. 1–15.
- Zhao et al. (2024b) Xiaopeng Zhao, Shen Wang, Zhenlin An, and Lei Yang. 2024b. Crowdsourced Geospatial Intelligence: Constructing 3D Urban Maps with Satellitic Radiance Fields. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 8, 3 (2024), 1–24.
- Zhong et al. (2022) Yihan Zhong, Weisong Wen, Hoi-Fung Ng, Xiwei Bai, and Li-Ta Hsu. 2022. Real-time factor graph optimization aided by graduated non-convexity based outlier mitigation for smartphone decimeter challenge. In Proceedings of the 35th International Technical Meeting of the Satellite Division of The Institute of Navigation (ION GNSS+ 2022). 2339–2348.

### v2

@@ Acknowledgements.
- Wang et al. (2015b) Yang Wang, Rong Yang, Keck Voon Ling, and Eng Kee Poh. 2015b. Robust Vector Tracking Loop Using Moving Horizon Estimation. In Proceedings of the ION 2015 Pacific PNT Meeting. Institute of Navigation, Honolulu, Hawaii, 640–648.
- Wen et al. (2021) Weisong Wen, Tim Pfeifer, Xiwei Bai, and Li-Ta Hsu. 2021. Factor Graph Optimization for GNSS/INS Integration: A Comparison with the Extended Kalman Filter. NAVIGATION: Journal of the Institute of Navigation 68, 2 (2021), 315–331. doi:10.1002/navi.421
- Wen et al. (2019) Weisong Wen, Guohao Zhang, and Li-ta Hsu. 2019. Correcting NLOS by 3D LiDAR and Building Height to Improve GNSS Single Point Positioning. NAVIGATION: Journal of the Institute of Navigation 66, 4 (2019), 705–718. doi:10.1002/navi.335
- Weng et al. (2023) Duojie Weng, Zhiyu Hou, Yang Meng, Miaomiao Cai, and Yanyiu Chan. 2023. Characterization and Mitigation of Urban GNSS Multipath Effects on Smartphones. Measurement 223 (2023), 113766. doi:10.1016/j.measurement.2023.113766
- Weng et al. (2024a) Xu Weng, Yuhui Jin, and KV Ling. 2024a. GnssQuest: Questing for Suitable GNSS Satellites through Augmented Reality. In Proceedings of the 22nd ACM Conference on Embedded Networked Sensor Systems (Hangzhou, China). ACM, New York, NY, USA, 867–868. doi:10.1145/3666025.3699411
- Weng et al. (2024c) Xu Weng, K.V. Ling, Haochen Liu, and Kun Cao. 2024c. Towards End-to-End GPS Localization with Neural Pseudorange Correction. In 2024 27th International Conference on Information Fusion (FUSION). IEEE, Venice, Italy, 1–7. doi:10.23919/FUSION59988.2024.10706359
- Weng and Ling (2023) Xu Weng and Keck Voon Ling. 2023. Localization with Noisy Android Raw GNSS Measurements. In 2023 IEEE Asia Pacific Conference on Wireless and Mobile (APWiMob). IEEE, Bali, Indonesia, 95–101. doi:10.1109/APWiMob59963.2023.10365597
@@
- Xia et al. (2024) Xiao Xia, Weisong Wen, and Li-Ta Hsu. 2024. Integrity-Constrained Factor Graph Optimization for GNSS Positioning in Urban Canyons. NAVIGATION: Journal of the Institute of Navigation 71, 3 (2024), 1–27. doi:10.33012/navi.660
@@
- Xu et al. (2023b) Huatao Xu, Pengfei Zhou, Rui Tan, and Mo Li. 2023b. Practically Adopting Human Activity Recognition. In Proceedings of the 29th Annual International Conference on Mobile Computing and Networking (Madrid, Spain). ACM, New York, NY, USA, 1285–1299. doi:10.1145/3570361.3613299
- Xu and Hsu (2024) Penghui Xu and Li-Ta Hsu. 2024. AutoW: Self-Supervision Learning for Weighting Estimation in GNSS Positioning. In Proceedings of the 37th International Technical Meeting of the Satellite Division of The Institute of Navigation (ION GNSS+ 2024). ION, Baltimore, Maryland, 2630–2644. doi:10.33012/2024.19896
- Xu et al. (2023a) Penghui Xu, Hoi-Fung Ng, Yihan Zhong, Guohao Zhang, Weisong Wen, Bo Yang, and Li-Ta Hsu. 2023a. Differentiable Factor Graph Optimization with Intelligent Covariance Adaptation for Accurate Smartphone Positioning. In Proceedings of the 36th International Technical Meeting of the Satellite Division of The Institute of Navigation (ION GNSS+ 2023). ION, Denver, Colorado, 2765–2773. doi:10.33012/2023.19297
@@
- Yang et al. (2023) Fan Yang, Chen Wang, Cesar Cadena, and Marco Hutter. 2023. iPlanner: Imperative Path Planning. In Proceedings of Robotics: Science and Systems XIX. RSS Foundation, Daegu, Republic of Korea, 9 pages. doi:10.15607/RSS.2023.XIX.064
- Zhan et al. (2024) Zitong Zhan, Dasong Gao, Yun-Jou Lin, Youjie Xia, and Chen Wang. 2024. iMatching: Imperative Correspondence Learning. In Computer Vision – ECCV 2024. Springer Nature Switzerland, Cham, 183–200. doi:10.1007/978-3-031-72933-1_11
- Zhang et al. (2023) Aston Zhang, Zachary C Lipton, Mu Li, and Alexander J Smola. 2023. Dive into deep learning. Cambridge University Press, Shaftesbury Road Cambridge CB2 8EA, UK.
@@
- Zhao et al. (2025) Ling Zhao, Zhengliang Lu, Keck Voon Ling, Yuandong Hu, Kan Zheng, and Wenhe Liao. 2025. Multirate Cascade Spacecraft Attitude-Orbit Integrated State Estimation and Control Framework Based on MHE and PWA-MPC. IEEE Trans. Aerospace Electron. Systems 61, 4 (2025), 9990–10007. doi:10.1109/TAES.2025.3552750
- Zhao et al. (2023) Xiaopeng Zhao, Zhenlin An, Qingrui Pan, and Lei Yang. 2023. NeRF2: Neural Radio-Frequency Radiance Fields. In Proceedings of the 29th Annual International Conference on Mobile Computing and Networking (Madrid, Spain) (ACM MobiCom ’23). Association for Computing Machinery, New York, NY, USA, Article 27, 15 pages. doi:10.1145/3570361.3592527
- Zhao et al. (2024b) Xiaopeng Zhao, Shen Wang, Zhenlin An, and Lei Yang. 2024b. Crowdsourced Geospatial Intelligence: Constructing 3D Urban Maps with Satellitic Radiance Fields. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 8, 3, Article 146 (Sept. 2024), 24 pages. doi:10.1145/3678572
- Zhong et al. (2022) Yihan Zhong, Weisong Wen, Hoi-Fung Ng, Xiwei Bai, and Li-Ta Hsu. 2022. Real-time Factor Graph Optimization Aided by Graduated Non-convexity Based Outlier Mitigation for Smartphone Decimeter Challenge. In Proceedings of the 35th International Technical Meeting of the Satellite Division of The Institute of Navigation (ION GNSS+ 2022). ION, Denver, Colorado, 2339–2348. doi:10.33012/2022.18382